CAN YOU HELP ME SOLVE THESE EQUATIONS:
X+Y =9
X-Y =3
THANKS

Answer by ptaylor(2198)


X+Y=9---------------------------Equation #1
X-Y=3----------------------------Equation #2
First we will add Eq1 and Eq2 and we get
X+X+Y-Y=9+3 collect like terms
2X=12 divide both sides by 2
X=6 substitute 6 for X in Eq1 and we get:
6+Y=9 subtract 6 from both sides
6-6+Y=9-6 collect like terms
Y=3
So are answer is:
X=6, Y=3
CK
Substitute the values of X and Y into eq 1
6+3=9
9=9
Now substitute the values of X and Y into eq 2
6-3=3
3=3
Hope this helps----ptaylor